40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Sat Sep 6 12:54:05 2014, Tara, Evan, DailyProgress, optic, North photothermal TF nPT.pdfnpttf.zip
    Reply  Sun Sep 7 19:33:40 2014, Tara, Evan, DailyProgress, optic, North photothermal TF photothermalTF.pdfnorthPT.zip
       Reply  Sun Nov 16 23:32:49 2014, Evan, DailyProgress, optic, North photothermal TF 
Message ID: 1505     Entry time: Sun Sep 7 19:33:40 2014     In reply to: 1503     Reply to this: 1532
Author: Tara, Evan 
Type: DailyProgress 
Category: optic 
Subject: North photothermal TF 

Tara and I took another photothermal TF of the north cavity today. Relevant parameters:

  • Power incident on cavity: 10 mW (up from the usual 1 mW)
  • Beat frequency: 1.2 MHz, drifting to 650 kHz (we are hoping it will swing through 0 Hz overnight and settle above a few megahertz by tomorrow)
  • DC voltage on north ISS PD: 2.39(5) V
  • DC power transmitted through cavity: 3.77(2) mW
  • PLL actuation coefficient: 50 kHzpk / 1 Vrms
  • PLL UGF: 80 kHz (measured)
  • EOAM drive: 5 Vpp from 20 kHz to 300 Hz, then 3 Vpp from 300 Hz to 0.2 Hz

In the attached data I have already converted the raw data (in V/V) into hertz of beat frequency per watt of circulating power. For this I use the conversion factor (50 kHz / 21/2 V) × (2.39 V / 3.77 mW) × π / F, with F = 16 700. Since the TF (again) appears to be junk above 1 kHz, I haven't bothered undoing the CLTF of the PLL.

The attached plot shows the expected photothermal TF in terms of hertz of beat frequency per watt of absorbed power per mirror. Therefore, the scaling factor that makes our measurement (given in hertz per watt of circulating power) overlap with the expected TF (given in hertz per watt of absorbed power per mirror) should be the average absorption of each mirror. I find that this scaling factor is 6 ppm, which seems surprisingly low, especially given our earlier finding that we have at least 120 ppm of scatter + absorption loss. So I will double check for missing factors of 2, 4, π, etc.

At any rate, the shape of the measured transfer function appears to be in good agreement with the expectation up to 100 Hz. If we believe that the coating/substrate photothermal crossover happens around 10 Hz, and we believe our measurement from 10 Hz to 100 Hz, then this seems to indicate that the thermo-optic cancellation has been somewhat successful.

Attachment 1: photothermalTF.pdf  102 kB  Uploaded Sun Sep 7 20:49:01 2014  | Hide | Hide all
Attachment 2: northPT.zip  106 kB  Uploaded Sun Sep 7 20:49:32 2014
ELOG V3.1.3-